# EMNLP 2024 Paraphrase Types Elicit Prompt Engineering Capabilities

### v4

@@ Abstract
Much of the success of modern language models depends on finding a suitable prompt to instruct the model. Until now, it has been largely unknown how variations in the linguistic expression of prompts affect these models. This study systematically and empirically evaluates which linguistic features influence models through paraphrase types, i.e., different linguistic changes at particular positions. We measure behavioral changes for five models across 120 tasks and six families of paraphrases (i.e., morphology, syntax, lexicon, lexico-syntax, discourse, and others). We also control for other prompt engineering factors (e.g., prompt length, lexical diversity, and proximity to training data). Our results show a potential for language models to improve tasks when their prompts are adapted in specific paraphrase types (e.g., 6.7% median gain in Mixtral 8x7B; 5.5% in LLaMA 3 8B; cf. Figure 1). In particular, changes in morphology and lexicon, i.e., the vocabulary used, showed promise in improving prompts. These findings contribute to developing more robust language models capable of handling variability in linguistic expression. Code: https://github.com/jpwahle/emnlp24-prompt-paraphrase.
@@ 1 Introduction
So far, there has been little work to systematically evaluate how linguistic variations of prompts affect current language models. We fill this gap by evaluating hundreds of thousands of prompt variations across 120 tasks and five models (3.24m total prompt-example combinations). Our procedure is outlined in Figure 2. We measure how sensitive the models are to different prompt changes and what we need to improve their robustness and effectiveness in performing different tasks. Our work addresses the following key questions:
@@
Our results show that adapting prompts to specific types (e.g., morphology) can yield significant gains in many models (cf. Figure 1) across different downstream tasks, such as summarization, text completion, or dialogue generation. We also control for various confounding factors, such as the length of prompts, lexical diversity, and proximity to training data, and show that performance gains can be observed independent of these additional factors. We recommend which paraphrase types to consider when adapting prompts for a particular task (e.g., polarity substitutions for sentiment analysis).
@@ 3 Data & Models
Central to this study is a dataset of tasks, prompts, and variations of these prompts. We use the Super-NaturalInstructions dataset Wang et al. (2022b), which contains more than 1 600 tasks with their respective prompts. We describe in detail how we construct different variations of these prompts in the experiments (Section 4). We sample 120 different tasks (each task has its own dataset) from 24 different task families (e.g., question answering, sentiment analysis) with the following conditions: the task family must have at least 10 tasks, the primary language is English, and each task has 200 examples to provide sufficient statistical power. We sample 5 tasks per domain, leaving us with 24 task families x 5 tasks per domain = 120 total datasets. See Appendix A for the task details.
We choose the five best non-proprietary models according to the LMSYS chatbot arena leaderboard in descending order of performance as of 1 May 2024111https://chat.lmsys.org/?leaderboard: LLaMA 3 Instruct (70B), Command R+ (104B), Mixtral 8x7B Instruct (47B), LLaMA 3 Instruct (8B), and Gemma Instruct (7B). Unless otherwise noted, we set the temperature to 0.2, the probability mass to sample words (top p) to 0.9, the penalty for repeatedly sampling the same sequence of tokens to 0.1, and the maximum number of tokens generated to the average of human references for that task. This number varies for different tasks but averages between 1 and 3154 tokens (see Appendix A for details). We use 40 NVIDIA A100 GPUs (40GB) with 16-bit precision (8 A100s for each model). Our experiments required a total computational budget of approximately 2880 A100 GPU hours, resulting in 311kg of CO2 equivalent [^2] https://mlco2.github.io/impact/#compute.
@@ 4 Experiments
Ans. We start with a seed prompt designed for a specific task, e.g., “In this task, you are given Twitter posts. Your task is to label the post’s emotion as sadness, joy, disgust, anger, fear, or surprise.” We paraphrase each prompt with one of 26 linguistic changes (e.g., syntax, lexicon) using the gpt-3.5-turbo-16k model, which we fine-tuned in the same way as Wahle et al. (2023) using the ETPC dataset Kovatchev et al. (2018). See Table 9 and Figures 8 and 10 in Appendix A for an extensive list of types, full prompt template and examples respectively. For the original prompt and its 26 paraphrases, we test each model’s few-shot settings. Depending on the task, the number of few-shot examples ranges from 3 to 6 (sum of positive and negative examples).
@@
Results. As Figure 1 already revealed, we can observe marked changes when adjusting prompts. Command R+ experiences up to 6.3% median performance gain, and Gemma 7B up to 13.4%. We decompose these results by the different types in Figure 3. Lexicon changes (+1.26%), closely followed by syntax changes (+1.19%), account for the largest median performance gain. The potential loss is comparable across changes. Overall, paraphrases had more upside than downside potential (+1.16% vs. -0.87% median change).
@@
Results. Figure 4 shows that title generation (+6.01%), text completion (+5.86%), and question answering (+5.60%) gain the most performance while having a low potential for a negative impact. Commonsense classification (-4.86%), sentiment analysis (-4.83%), and word semantics (-4.82%) have the highest loss potential.
Specific perturbations of the prompt affect the performance of models on tasks from varying domains differently. Morphology changes in the prompt have the largest gains in wrong candidate generation (+26.0%), question generation (+21.5%), and textual entailment (+17.5%). Lexical changes show consistent gains in summarization (10.8%), wrong candidate generation (+8.9%), and title generation (+7.1%). These tasks rely on semantic precision and vocabulary richness to interpret or generate nuanced responses. Positive effects with discourse changes can be observed in tasks that require an understanding of longer contexts or multi-sentence structures, such as summarization (+7.8%). Detailed results for each model are available in Tables 10, 12, 13, 11, 15 and 14 in Appendix C.
Discussion. Some tasks are more sensitive to instructions than others. Tasks such as question generation and text completion may benefit from syntax changes because they can improve grammatical accuracy while reducing the flexibility of more creative tasks such as dialog generation. In particular, sentiment classification seems to benefit from polarity substitutions and is sensitive to negation, consistent with related work Wiegand et al. (2010). Lexical changes improve performance in vocabulary-intensive tasks such as named entity recognition and text entailment. However, they may hinder performance in more general contexts such as question rewriting, which is often based on data from Wikipedia and knowledge bases (see Table 5 in Appendix A). In discourse, better context management, as well as indirect and direct style, improves clarity and coherence in tasks like summarization and long-form answer generation.
@@
Results. In Figure 1, note how the smaller 8B version of LLaMA 3 can gain up to 5.5% median task performance, while the larger 70B model can gain much less with up to 2.8%. LLaMA 3 8B has a lower baseline performance of 0.58, while the 70B model’s baseline is at 0.65 (see Figure 12 in Appendix C for more details).
@@
If we always choose the best (paraphrased) prompts, we can achieve even higher performance gains; for example, LLaMA 3 8B could gain 21.1%, making the model markedly better than its 70B counterpart by tweaking the prompts. We want to note that it is difficult to always find the best prompt. Therefore, we report median performance gains in the main body of the paper and report results for selecting the best prompt in Figure 9 in Appendix C.
@@
Results. The x-axis of Figure 5 shows $\Delta_{train}$, which represents this difference in similarity to the training data between the paraphrased and original prompts. The y-axis then shows downstream task performance. The results show few very close training examples for our tasks, i.e., the ROUGE-L between prompt and closest FineWeb example is typically below 0.5. Most successful paraphrased prompts with a downstream task performance greater than 0.8 do not show higher similarity to training than the original prompt. Sometimes, higher scores can be achieved when paraphrases are closer to the training (top right of the figure).
@@
We measure lexical richness with four metrics that focus on the diversity of tokens and text segments Guiraud (1960); Maas (1972); Covington (2007); McCarthy and Jarvis (2010): Root Type-Token Ratio (RTTR), Maas, Moving Average Type-Token Ratio (MATTR), and Measure of Lexical Diversity (MTLD) — more details in Appendix B. We select four task families with long responses to measure lexical richness: summarization, question generation, question rewriting, and dialogue generation (20 tasks in total).
Results. Summarization tasks have a particularly high diversity in responses as measured by Maas, RTTR, and MTLD (Figure 6 (a), (c), (d)). Overall, lexicon and lexico-syntactic changes often lead to higher performance at the expense of lexical richness. Models produce responses with lower lexical diversity in question rewriting tasks as measured by MAAS (Figure 6 (a)). Syntax and other changes lead to both lower lexical diversity and task performance gains compared to others measured by RTTR (Figure 6 (c)). Question rewriting achieves the highest lexical diversity as measured by MAAS, specifically with lexicon and lexico-syntactic changes (Figure 6 (a)), and question generation and question rewriting have overall lower diversity and performance gains, specifically with discourse and other changes (Figure 6 (c)).
@@
Ans. We use three markers to describe the complexity of a prompt paraphrase relative to the original instruction: the deviation in absolute number of tokens, word position deviation, and lexical deviation Liu and Soh (2022). We use Pearson correlation to measure a quantitative relationship between these three markers of prompt complexity and task performance. More details about the metrics and correlations can be found in Appendix B.
@@
Results. Across tasks, there is no significant correlation between changes in prompt length, changes in word position, and lexical deviation, as all correlations are close to zero (see Table 2 in the Appendix C). We found no evidence that making prompts more or less complex is associated with higher or lower task performance.
@@
Results. Figure 7 shows contour plots of different temperature choices. The coloring shows $\Delta_{temp}$ for different temperatures for the paraphrased (x-axis) and original (y-axis) prompts. Note how the distributions within the summarization have similar hills and valleys (see Figure 7 (c) and (d)). For title generation, with generally higher performance loss, the highest gains are achieved for low to medium temperatures (up to 0.6 for the paraphrased prompt). For summarization, with overall higher scores, low temperature yields the highest performance difference, close to 0 for the paraphrased prompts and between 0.2 and 0.4 for the original.
@@ Appendix A Details on the Prompts & Tasks
Prompts. We use the template shown in Figure 8 to construct prompts for each of the 120 tasks. In Figure 10, we show an example for the numerical commonsense reasoning task NumerSense in which the model has to predict a blank token ‘_’ Lin et al. (2020). To paraphrase prompts, we use one of 26 paraphrase types in six groups shown in Table 9. Five examples of paraphrases applied to prompts can be seen in Table 1. In the first case, the added specification made the instruction more precise; in the second case, the model removed seemingly unnecessary context. The third case involved replacing words or phrases with synonyms or near-synonyms that have the same meaning in the given context, while the fourth case involved restructuring the sentence or changing the way the information is presented while maintaining the same overall meaning. Finally, the fifth case replaced a single word with a phrase (or vice versa) that conveys the same meaning. Further paraphrase type definitions and perturbation examples can be found in Vila et al. (2014); Meier et al. (2024).
Tasks. We provide an overview of all 120 tasks together with the paraphrase type that, when applied to prompts in that task domain, had the most positive influence together with the potential downstream task performance gain over the original prompt in Tables 4, 5, 6 and 7. Further, we want to note an observation of our experiments showing that tasks containing content judged as unethical (such as toxic language detection on Twitter) led to many denied responses by the chat or instruction-finetuned models. Likely because of their post-training, models answered that they could not respond to these questions, although we did not ask them to produce new harmful content but to classify existing content. We removed these tasks from our final analysis. This observation raises questions about how many powerful models can be used in the future to classify toxic content when their guardrails prevent them from answering.
@@
Figure 9: The maximum performance gain models can experience if their prompts are adjusted in optimal ways as explained in Table 8.
@@
Output: 4   # Not expressed in words but numbers. Input: Numbers less than _ are called negative. Output: one   # Logically wrong. Input: Some plant varieties can grow up to _ feet tall. Output:
@@ B.3 Lexical Diversity Metrics
MTLD=Total tokensNumber of $\text{MTLD}=\frac{\text{Total tokens}}{\text{Number of segments}}$
@@ B.5 Potential Downstream Gain
Figure 1 has shown the potential median gain that can be reached by paraphrasing prompts in specific types. In Table 8, we show how the median is calculated per example. For each example in a task dataset, we take the downstream task scores of paraphrased prompts that are higher than the original prompt and take the median. Then, for all examples in the dataset, we average that median.
Further, one can also compute the maximum performance that can be reached if the best paraphrase type would be chosen to paraphrase the prompt. As previous experiments have shown, it is challenging to devise a robust method to select the best paraphrase of a prompt. More research on the Monte-Carlo approach could bring us closer to this maximum. The median is a fair way to represent a gain that users could expect by selecting one of the multiple paraphrases of prompts that are better than the original. Compared to the mean, the median is not influenced by outliers that are unlikely to select (e.g., the two 1.0 scores in Table 8 are unlikely to reach; the most likely expectation is 0.67, which is also the median).

### v5

@@ Abstract
Much of the success of modern language models depends on finding a suitable prompt to instruct the model. Until now, it has been largely unknown how variations in the linguistic expression of prompts affect these models. This study systematically and empirically evaluates which linguistic features influence models through paraphrase types, i.e., different linguistic changes at particular positions. We measure behavioral changes for five models across 120 tasks and six families of paraphrases (i.e., morphology, syntax, lexicon, lexico-syntax, discourse, and others). We also control for other prompt engineering factors (e.g., prompt length, lexical diversity, and proximity to training data). Our results show a potential for language models to improve tasks when their prompts are adapted in specific paraphrase types (e.g., 6.7% median gain in Mixtral 8x7B; 5.5% in LLaMA 3 8B; cf. Figure˜1). In particular, changes in morphology and lexicon, i.e., the vocabulary used, showed promise in improving prompts. These findings contribute to developing more robust language models capable of handling variability in linguistic expression. Code: https://github.com/jpwahle/emnlp24-prompt-paraphrase.
@@ 1 Introduction
So far, there has been little work to systematically evaluate how linguistic variations of prompts affect current language models. We fill this gap by evaluating hundreds of thousands of prompt variations across 120 tasks and five models (3.24m total prompt-example combinations). Our procedure is outlined in Figure˜2. We measure how sensitive the models are to different prompt changes and what we need to improve their robustness and effectiveness in performing different tasks. Our work addresses the following key questions:
@@
Our results show that adapting prompts to specific types (e.g., morphology) can yield significant gains in many models (cf. Figure˜1) across different downstream tasks, such as summarization, text completion, or dialogue generation. We also control for various confounding factors, such as the length of prompts, lexical diversity, and proximity to training data, and show that performance gains can be observed independent of these additional factors. We recommend which paraphrase types to consider when adapting prompts for a particular task (e.g., polarity substitutions for sentiment analysis).
@@ 3 Data & Models
Central to this study is a dataset of tasks, prompts, and variations of these prompts. We use the Super-NaturalInstructions dataset Wang et al. (2022b), which contains more than 1 600 tasks with their respective prompts. We describe in detail how we construct different variations of these prompts in the experiments (Section˜4). We sample 120 different tasks (each task has its own dataset) from 24 different task families (e.g., question answering, sentiment analysis) with the following conditions: the task family must have at least 10 tasks, the primary language is English, and each task has 200 examples to provide sufficient statistical power. We sample 5 tasks per domain, leaving us with 24 task families x 5 tasks per domain = 120 total datasets. See Appendix˜A for the task details.
We choose the five best non-proprietary models according to the LMSYS chatbot arena leaderboard in descending order of performance as of 1 May 2024111https://chat.lmsys.org/?leaderboard: LLaMA 3 Instruct (70B), Command R+ (104B), Mixtral 8x7B Instruct (47B), LLaMA 3 Instruct (8B), and Gemma Instruct (7B). Unless otherwise noted, we set the temperature to 0.2, the probability mass to sample words (top p) to 0.9, the penalty for repeatedly sampling the same sequence of tokens to 0.1, and the maximum number of tokens generated to the average of human references for that task. This number varies for different tasks but averages between 1 and 3154 tokens (see Appendix˜A for details). We use 40 NVIDIA A100 GPUs (40GB) with 16-bit precision (8 A100s for each model). Our experiments required a total computational budget of approximately 2880 A100 GPU hours, resulting in 311kg of CO2 equivalent [^2] https://mlco2.github.io/impact/#compute.
@@ 4 Experiments
Ans. We start with a seed prompt designed for a specific task, e.g., “In this task, you are given Twitter posts. Your task is to label the post’s emotion as sadness, joy, disgust, anger, fear, or surprise.” We paraphrase each prompt with one of 26 linguistic changes (e.g., syntax, lexicon) using the gpt-3.5-turbo-16k model, which we fine-tuned in the same way as Wahle et al. (2023) using the ETPC dataset Kovatchev et al. (2018). See Table˜9 and Figures˜8 and 10 in Appendix˜A for an extensive list of types, full prompt template and examples respectively. For the original prompt and its 26 paraphrases, we test each model’s few-shot settings. Depending on the task, the number of few-shot examples ranges from 3 to 6 (sum of positive and negative examples).
@@
Results. As Figure˜1 already revealed, we can observe marked changes when adjusting prompts. Command R+ experiences up to 6.3% median performance gain, and Gemma 7B up to 13.4%. We decompose these results by the different types in Figure˜3. Lexicon changes (+1.26%), closely followed by syntax changes (+1.19%), account for the largest median performance gain. The potential loss is comparable across changes. Overall, paraphrases had more upside than downside potential (+1.16% vs. -0.87% median change).
@@
Results. Figure˜4 shows that title generation (+6.01%), text completion (+5.86%), and question answering (+5.60%) gain the most performance while having a low potential for a negative impact. Commonsense classification (-4.86%), sentiment analysis (-4.83%), and word semantics (-4.82%) have the highest loss potential.
Specific perturbations of the prompt affect the performance of models on tasks from varying domains differently. Morphology changes in the prompt have the largest gains in wrong candidate generation (+26.0%), question generation (+21.5%), and textual entailment (+17.5%). Lexical changes show consistent gains in summarization (10.8%), wrong candidate generation (+8.9%), and title generation (+7.1%). These tasks rely on semantic precision and vocabulary richness to interpret or generate nuanced responses. Positive effects with discourse changes can be observed in tasks that require an understanding of longer contexts or multi-sentence structures, such as summarization (+7.8%). Detailed results for each model are available in Tables˜10, 12, 13, 11, 15 and 14 in Appendix˜C.
Discussion. Some tasks are more sensitive to instructions than others. Tasks such as question generation and text completion may benefit from syntax changes because they can improve grammatical accuracy while reducing the flexibility of more creative tasks such as dialog generation. In particular, sentiment classification seems to benefit from polarity substitutions and is sensitive to negation, consistent with related work Wiegand et al. (2010). Lexical changes improve performance in vocabulary-intensive tasks such as named entity recognition and text entailment. However, they may hinder performance in more general contexts such as question rewriting, which is often based on data from Wikipedia and knowledge bases (see Table˜5 in Appendix˜A). In discourse, better context management, as well as indirect and direct style, improves clarity and coherence in tasks like summarization and long-form answer generation.
@@
Results. In Figure˜1, note how the smaller 8B version of LLaMA 3 can gain up to 5.5% median task performance, while the larger 70B model can gain much less with up to 2.8%. LLaMA 3 8B has a lower baseline performance of 0.58, while the 70B model’s baseline is at 0.65 (see Figure˜12 in Appendix˜C for more details).
@@
If we always choose the best (paraphrased) prompts, we can achieve even higher performance gains; for example, LLaMA 3 8B could gain 21.1%, making the model markedly better than its 70B counterpart by tweaking the prompts. We want to note that it is difficult to always find the best prompt. Therefore, we report median performance gains in the main body of the paper and report results for selecting the best prompt in Figure˜9 in Appendix˜C.
@@
Results. The x-axis of Figure˜5 shows $\Delta_{train}$, which represents this difference in similarity to the training data between the paraphrased and original prompts. The y-axis then shows downstream task performance. The results show few very close training examples for our tasks, i.e., the ROUGE-L between prompt and closest FineWeb example is typically below 0.5. Most successful paraphrased prompts with a downstream task performance greater than 0.8 do not show higher similarity to training than the original prompt. Sometimes, higher scores can be achieved when paraphrases are closer to the training (top right of the figure).
@@
We measure lexical richness with four metrics that focus on the diversity of tokens and text segments Guiraud (1960); Maas (1972); Covington (2007); McCarthy and Jarvis (2010): Root Type-Token Ratio (RTTR), Maas, Moving Average Type-Token Ratio (MATTR), and Measure of Lexical Diversity (MTLD) — more details in Appendix˜B. We select four task families with long responses to measure lexical richness: summarization, question generation, question rewriting, and dialogue generation (20 tasks in total).
Results. Summarization tasks have a particularly high diversity in responses as measured by Maas, RTTR, and MTLD (Figure˜6 (a), (c), (d)). Overall, lexicon and lexico-syntactic changes often lead to higher performance at the expense of lexical richness. Models produce responses with lower lexical diversity in question rewriting tasks as measured by MAAS (Figure˜6 (a)). Syntax and other changes lead to both lower lexical diversity and task performance gains compared to others measured by RTTR (Figure˜6 (c)). Question rewriting achieves the highest lexical diversity as measured by MAAS, specifically with lexicon and lexico-syntactic changes (Figure˜6 (a)), and question generation and question rewriting have overall lower diversity and performance gains, specifically with discourse and other changes (Figure˜6 (c)).
@@
Ans. We use three markers to describe the complexity of a prompt paraphrase relative to the original instruction: the deviation in absolute number of tokens, word position deviation, and lexical deviation Liu and Soh (2022). We use Pearson correlation to measure a quantitative relationship between these three markers of prompt complexity and task performance. More details about the metrics and correlations can be found in Appendix˜B.
@@
Results. Across tasks, there is no significant correlation between changes in prompt length, changes in word position, and lexical deviation, as all correlations are close to zero (see Table˜2 in the Appendix˜C). We found no evidence that making prompts more or less complex is associated with higher or lower task performance.
@@
Results. Figure˜7 shows contour plots of different temperature choices. The coloring shows $\Delta_{temp}$ for different temperatures for the paraphrased (x-axis) and original (y-axis) prompts. Note how the distributions within the summarization have similar hills and valleys (see Figure˜7 (c) and (d)). For title generation, with generally higher performance loss, the highest gains are achieved for low to medium temperatures (up to 0.6 for the paraphrased prompt). For summarization, with overall higher scores, low temperature yields the highest performance difference, close to 0 for the paraphrased prompts and between 0.2 and 0.4 for the original.
@@ Appendix A Details on the Prompts & Tasks
Prompts. We use the template shown in Figure˜8 to construct prompts for each of the 120 tasks. In Figure˜10, we show an example for the numerical commonsense reasoning task NumerSense in which the model has to predict a blank token ‘_’ Lin et al. (2020). To paraphrase prompts, we use one of 26 paraphrase types in six groups shown in Table˜9. Five examples of paraphrases applied to prompts can be seen in Table˜1. In the first case, the added specification made the instruction more precise; in the second case, the model removed seemingly unnecessary context. The third case involved replacing words or phrases with synonyms or near-synonyms that have the same meaning in the given context, while the fourth case involved restructuring the sentence or changing the way the information is presented while maintaining the same overall meaning. Finally, the fifth case replaced a single word with a phrase (or vice versa) that conveys the same meaning. Further paraphrase type definitions and perturbation examples can be found in Vila et al. (2014); Meier et al. (2025).
Tasks. We provide an overview of all 120 tasks together with the paraphrase type that, when applied to prompts in that task domain, had the most positive influence together with the potential downstream task performance gain over the original prompt in Tables˜4, 5, 6 and 7. Further, we want to note an observation of our experiments showing that tasks containing content judged as unethical (such as toxic language detection on Twitter) led to many denied responses by the chat or instruction-finetuned models. Likely because of their post-training, models answered that they could not respond to these questions, although we did not ask them to produce new harmful content but to classify existing content. We removed these tasks from our final analysis. This observation raises questions about how many powerful models can be used in the future to classify toxic content when their guardrails prevent them from answering.
@@
Figure 9: The maximum performance gain models can experience if their prompts are adjusted in optimal ways as explained in Table˜8.
@@
Output: 4   # Not expressed in words but numbers.
     Input: Numbers less than _ are called negative.
Output: one   # Logically wrong.
Input: Some plant varieties can grow up to _ feet tall. Output:
@@ B.3 Lexical Diversity Metrics
$\text{MTLD}=\frac{\text{Total tokens}}{\text{Number of segments}}$
@@ B.5 Potential Downstream Gain
Figure˜1 has shown the potential median gain that can be reached by paraphrasing prompts in specific types. In Table˜8, we show how the median is calculated per example. For each example in a task dataset, we take the downstream task scores of paraphrased prompts that are higher than the original prompt and take the median. Then, for all examples in the dataset, we average that median.
Further, one can also compute the maximum performance that can be reached if the best paraphrase type would be chosen to paraphrase the prompt. As previous experiments have shown, it is challenging to devise a robust method to select the best paraphrase of a prompt. More research on the Monte-Carlo approach could bring us closer to this maximum. The median is a fair way to represent a gain that users could expect by selecting one of the multiple paraphrases of prompts that are better than the original. Compared to the mean, the median is not influenced by outliers that are unlikely to select (e.g., the two 1.0 scores in Table˜8 are unlikely to reach; the most likely expectation is 0.67, which is also the median).